OpenStudy (anonymous):

In simplest form, what is the ratio of 3 feet to 18 inches?

OpenStudy (anonymous):

\[\frac{3}{18}=\frac{1}{6}\]

OpenStudy (tkhunny):

\[\frac{ 3' }{ 18" } = \frac{ 36" }{ 18" } = \frac{ 36 }{ 18 } = 2\] The units need to match up!